OpenStudy (anonymous):

Which placement of grouping symbols in the expression gives it a value of 22? 3 • 4 – 2 + 5 +7 A. 3 • [4 – (2 + 5) + 7] B. (3 • 4) – [2 + (5 + 7)] C. [(3 • 4) – 2] + (5 + 7) D. 3 • (4 – 2) + (5 + 7)

OpenStudy (anonymous):

@Adjax can you please help me?

OpenStudy (anonymous):

C. [(3 • 4) – 2] + (5 + 7)

OpenStudy (mokeira):

Choice C

OpenStudy (anonymous):

@wormie2014:: C: \[[(3 \times 4) -2] + (5+7) = [12 -2] + 12 = 10 + 12 = 22\] Remember that '.' = x(multiplication symbol)